# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ IV.2.2 $\overline{B}\to\overline{D}T$ decay amplitudes
Note that three out of four flavor exotic states, namely $T^{0}_{cs}$, $T^{+}_{cs}$ and $T^{0}_{c\bar{s}}$, are involved in these $\overline{B}\to\overline{D}T$ decays, while the other flavor exotic state, $T^{++}_{c\bar{s}}$ is absent. The absence of the doubly charged state in $\overline{B}\to\overline{D}T$ decays can be easily understood from charge conservation. We need to consider both $\Delta S=-1$ and $\Delta S=0$ transitions in $\overline{B}\to\overline{D}T$ decays to have both $T^{0}_{cs}$ and $T^{0}_{c\bar{s}}$, but without $T^{++}_{c\bar{s}}$.
@@
We turn to scenario II. Likewise, from Table 10, the decay modes involving $T$ in $\overline{\bf 15}$ only have contributions from internal $W$ amplitudes $C_{\bar{D}T}$ or $C^{\prime}_{\bar{D}T}$, and, consequently, their amplitudes are highly related. In particular, the modes involving flavor exotic $T$ have the following relations on decay rates,
$\Gamma(\overline{B}{}^{0}\to D^{-}T^{+}_{cs})$ $=$ $2\Gamma(B^{-}\to D^{-}T^{0}_{cs})=2\Gamma(\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs})=\Gamma(B^{-}\to\overline{D}^{0}T^{-}_{cs})$ (47)
$=$ $2\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D_{s}^{-}T^{0}_{c\bar{s}})=2\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{0}_{c\bar{s}})$
$=$ $4\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D_{s}^{-}T^{+}_{c\bar{s}})=4\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to D^{-}T^{+}_{c\bar{s}})$
$=$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{+}_{c\bar{s}\bar{s}})=3\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D^{-}T^{+}_{c})$
$=$ $3\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D^{-}T^{0}_{c})=3\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{c})$
$=$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to\overline{D}^{0}T^{-}_{c}).$
@@
We have nine out of twelve flavor exotic states involved in these $\overline{B}\to\overline{D}T$ decays, while all doubly chared states, $T^{++}_{c\bar{s}}$, $T^{++}_{c\bar{s}\bar{s}}$ and $T^{++}_{c}$, are absent, as $\overline{B}\to\overline{D}T^{++}$ decays are forbiden by charge conservation.
$\overline{B}\to D\overline{T}$, $\Delta S=-1$checkmark $\checkmark$ X $C_{\bar{T}}$
$\overline{B}\to\overline{D}T$, $\Delta S=-1$ X X $\checkmark$ $C_{\bar{D}T}$
$\overline{B}\to D\overline{T}$, $\Delta S=0$checkmark X $\checkmark$ $C^{\prime}_{\bar{T}}$
$\overline{B}\to\overline{D}T$, $\Delta S=0$ X $\checkmark$ X $C^{\prime}_{\bar{D}T}$
@@
$\Gamma(\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c})$ $=$ $\Gamma(B^{-}\to D_{s}^{-}T^{\prime\prime 0}_{c})=\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D^{-}T^{\prime\prime 0}_{c})$ (49)
$=$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c\bar{s}}),$
$\Gamma(\overline{B}{}_{s}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}})=\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D^{-}T^{\prime\prime+}_{c}),$ (50)
@@
$\Gamma(\overline{B}{}_{s}^{0}\to D^{-}T^{\prime\prime+}_{c})$ $=$ $\Gamma(\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{\prime\prime 0}_{c})=\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to\overline{D}^{0}T^{\prime\prime 0}_{c})$ (51)
$=$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D_{s}^{-}T^{\prime\prime+}_{c\bar{s}}).$
As they have contributions from external $W$ amplitudes, $T_{\bar{D}}$ or $T^{\prime}_{\bar{D}}$, the rates in Eqs. (49) and (50) are expected to be larger than those in Eq. (51) in the same $\Delta S$ transition.
In ref. Qin:2022nof the amplitudes of the following four modes in $\Delta S=-1$ transitions in scenario II, namely $\overline{B}{}^{0}\to D^{-}T^{+}_{cs}$, $B^{-}\to D^{-}T^{0}_{cs}$, $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs}$ and $B^{-}\to\overline{D}^{0}T^{-}_{cs}$ decays, were also given. Their results agree with those shown in Table 10, but with $C_{\bar{D}T}$ replaced with $T_{3}V_{cb}V^{*}_{cs}$, while the rest in Table 10 and all in Table 9 are new.
Before we end this section, it will be useful to compare the participation of $T^{*}_{c\bar{s}0}(2900)^{++}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ in $\overline{B}\to D\overline{T}$ and $\overline{B}\to\overline{D}T$ decays in $\Delta S=-1$ and $\Delta S=0$ transitions. We show in Table 11 such a comparison with the relevant topological amplitudes listed. It is evident that using only $\Delta S=-1$ transitions, one is unable to check if these three states belong to the same multiplet or not, as $C_{\bar{T}}$ and $C_{\bar{D}T}$ are unrelated, unlike $C_{\bar{T}}$ and $C^{\prime}_{\bar{T}}$ or $C_{\bar{D}T}$ and $C^{\prime}_{\bar{D}T}$. Therefore, to verify that, in any case, we need to consider the CKM suppressed $\Delta S=0$ transitions as well.
@@ V.2.1 $\overline{B}\to TP$ decay amplitudes
The $\overline{B}\to TP$ decay amplitudes, decomposed in these topological amplitudes, are shown in Table 12 for $\Delta S=0$ and $-1$ transitions in scenario I, and in Tables 13 and 14 for $\Delta S=0$ and $\Delta S=-1$ transitions in scenario II, respectively. These results can be applied to $\overline{B}\to TV$ decays by replacing $P$ with $V$ suitably.
@@
$\Gamma(\overline{B}{}^{0}\to T^{0}_{c\bar{s}}\overline{K}{}^{0})$ $=$ $\left|\frac{V_{ud}}{V_{us}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to T^{0}_{cs}K^{0}),$ (59)
@@
# Mode $A(\overline{B}_{q}\to T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}P)$ Mode $A(\overline{B}_{q}\to T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}P)$
$2^{*}$ $\overline{B}{}^{0}\to T^{+}_{c\bar{s}}K^{-}$ $-\frac{1}{\sqrt{2}}(C_{TP}-E_{\bar{B}T})$ $2^{*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}\pi^{-}$ $\frac{1}{\sqrt{2}}(C_{TP}-C_{PT,1})$

### v3

@@ IV.2.2 $\overline{B}\to\overline{D}T$ decay amplitudes
where the approximation in the last line neglects the penguin contributions. Note that three out of four flavor exotic states, namely $T^{0}_{cs}$, $T^{+}_{cs}$ and $T^{0}_{c\bar{s}}$, are involved in these $\overline{B}\to\overline{D}T$ decays, while the other flavor exotic state, $T^{++}_{c\bar{s}}$ is absent. The absence of the doubly charged state in $\overline{B}\to\overline{D}T$ decays can be easily understood from charge conservation. We need to consider both $\Delta S=-1$ and $\Delta S=0$ transitions in $\overline{B}\to\overline{D}T$ decays to have both $T^{0}_{cs}$ and $T^{0}_{c\bar{s}}$ final states.
One can roughly estimate the error in the above approximation by using
$\frac{|P_{T}|}{|C_{\bar{D}T}|}\sim\left|\frac{V_{tb}V^{*}_{ts}(c_{4}+\frac{c_{3}}{3})}{V_{cb}V^{*}_{cs}(c_{2}+\frac{c_{1}}{3})}\right|\simeq 0.19,$ (70)
and
$\frac{|P^{\prime}_{T}|}{|C^{\prime}_{\bar{D}T}|}\sim\left|\frac{V_{tb}V^{*}_{td}(c_{4}+\frac{c_{3}}{3})}{V_{cb}V^{*}_{cd}(c_{2}+\frac{c_{1}}{3})}\right|\simeq 0.17,$ (71)
where $c_{i}$ are the Wilson coefficients, see Beneke:2001ev . Therefore, the errors of neglecting penguin contributions are roughly estimated to be of the order $\lambda$. To be more precise, we can similarly define the following ratio:
$\tilde{R} \displaystyle\equiv$ $\left|\frac{V_{cb}V^{*}_{cd}}{V_{cb}V^{*}_{cs}}\right|^{2}\frac{\Gamma_{5}}{\Gamma_{6}}$
$=$ $\left|\frac{V_{cb}V^{*}_{cd}}{V_{cb}V^{*}_{cs}}\right|^{2}\frac{|C_{\bar{D}T}+P_{T}|^{2}+|\bar{C}_{\bar{D}T}+\bar{P}_{T}|^{2}}{|C^{\prime}_{\bar{D}T}+P^{\prime}_{T}|^{2}+|\bar{C}^{\prime}_{\bar{D}T}+\bar{P}^{\prime}_{T}|^{2}}$
$=$ $\left|\frac{V_{cb}V^{*}_{cd}}{V_{cb}V^{*}_{cs}}\right|^{2}$
$\times\frac{|V_{cb}V^{*}_{cd}(c_{2}+\frac{c_{1}}{3})+\tilde{\rho}e^{i\tilde{\phi}}V_{tb}V^{*}_{ts}(c_{4}+\frac{c_{3}}{3})|^{2}+|V^{*}_{cb}V_{cd}(c_{2}+\frac{c_{1}}{3})+\tilde{\rho}e^{i\tilde{\phi}}V^{*}_{tb}V_{ts}(c_{4}+\frac{c_{3}}{3})|^{2}}{|V_{cb}V^{*}_{cd}(c_{2}+\frac{c_{1}}{3})+\tilde{\rho}e^{i\tilde{\phi}}V_{tb}V^{*}_{td}(c_{4}+\frac{c_{3}}{3})|^{2}+|V^{*}_{cb}V_{cd}(c_{2}+\frac{c_{1}}{3})+\tilde{\rho}e^{i\tilde{\phi}}V^{*}_{tb}V_{td}(c_{4}+\frac{c_{3}}{3})|^{2}},$
where we introduce a correction factor $\tilde{\rho}$ and a relative strong phase $\phi$ between the penguin and the tree amplitudes. Again by taking $\tilde{\rho}=0\sim 2$ and $\tilde{\phi}=0\sim 2\pi$, we obtain
$\tilde{R}-1=-0.20\sim 0.09.$ (73)
Hence, the correction on the above approximation is estimated to be about 10% to 20% in the presence of penguin contributions. To effectively control the penguin pollution, further work is needed; see, for example, a study in a somewhat similar situation Fleischer:2007zn , where information of direct CP violation is needed. With this information, the sizes of the penguin and tree amplitudes, as well as the relative strong phase, can be probed.
@@
We turn to scenario II. Likewise, from Table 10, the decay modes involving $T$ in $\overline{\bf 15}$ only have contributions from $(C_{\bar{D}T}+P_{T})$ or $(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$, and, consequently, their amplitudes are highly related. In particular, the modes involving flavor exotic $T$ have the following relations on decay rates,
$\Gamma_{7}$ $\equiv$ $\Gamma(\overline{B}{}^{0}\to D^{-}T^{+}_{cs})=2\Gamma(B^{-}\to D^{-}T^{0}_{cs})=2\Gamma(\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs})=\Gamma(B^{-}\to\overline{D}^{0}T^{-}_{cs})$
$\Gamma_{8}$ $\equiv$ $2\Gamma(B^{-}\to D_{s}^{-}T^{0}_{c\bar{s}})=2\Gamma(\overline{B}{}_{s}^{0}\to\overline{D}^{0}T^{0}_{c\bar{s}})$
[…]
$=$ $\Gamma(B^{-}\to D^{-}T^{0}_{c})=3\Gamma(\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{c})=\Gamma(B^{-}\to\overline{D}^{0}T^{-}_{c}),$
$\Gamma_{7}$ $\simeq$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma_{8},$ (74)
@@
where the approximation in the last line neglects the penguin contributions and the correction to the approximation is about 10% to 20%, see Eq. (73). We have nine out of twelve flavor exotic states involved in these $\overline{B}\to\overline{D}T$ decays, while all doubly charged states, $T^{++}_{c\bar{s}}$, $T^{++}_{c\bar{s}\bar{s}}$ and $T^{++}_{c}$, are absent, as $\overline{B}\to\overline{D}T^{++}$ decays are forbiden by charge conservation.
@@
Note that the $\overline{B}^{0}\to\overline{D}^{0}T^{0}_{cs}$ decay has not been observed yet. It will be useful to search for this mode and check the relations in Eqs. (69) and (74).
@@
[…]
@@
[…]
where the approximation is neglecting penguin contributions and the correction to the approximation is about 10% to 20%, see Eq. (73). As they have contributions from external $W$ amplitudes, $T_{\bar{D}}$ or $T^{\prime}_{\bar{D}}$, the rates in Eqs. (76) and (77) are expected to be larger than those in Eq. (78) in the same $\Delta S$ transition.
In ref. Qin:2022nof the amplitudes of the following four modes in $\Delta S=-1$ transitions in scenario II, namely $\overline{B}{}^{0}\to D^{-}T^{+}_{cs}$, $B^{-}\to D^{-}T^{0}_{cs}$, $\overline{B}{}^{0}\to\overline{D}^{0}T^{0}_{cs}$ and $B^{-}\to\overline{D}^{0}T^{-}_{cs}$ decays, were also given. Their results agree with those shown in Table 10, but with $C_{\bar{D}T}$ replaced with $T_{3}V_{cb}V^{*}_{cs}$ and neglecting $P_{T}$, as they do not consider penguin contributions. The rest in Table 10 and all in Table 9 are new. Note that ref. Qin:2022nof also used a different notation for the open-charmed tetraquark states, see Table 2 for a comparison. For example, in Table 2 of ref. Qin:2022nof , one finds
$A(B^{-}\to D^{-}T_{cs\bar{u}\bar{d}})=T_{3}V_{cb}V^{*}_{cs},$ (79)
which corresponds to
$A(B^{-}\to D^{-}T^{0}_{cs})=C_{\bar{D}T}+P_{T},$ (80)
in Table 10 but with $C_{\bar{D}T}$ replaced with $T_{3}V_{cb}V^{*}_{cs}$ and neglecting $P_{T}$.
The final state $T$ can further decay to $DP$ and $DS$, or they can connect to those states virtually. One can obtain their full amplitudes using results in the section and those in Sec. III.
$\overline{B}\to D\overline{T}$, $\Delta S=-1$ $\checkmark$ $\checkmark$ X $(C_{\bar{T}}+P_{\bar{T},1})$
$\overline{B}\to\overline{D}T$, $\Delta S=-1$ X X $\checkmark$ $(C_{\bar{D}T}+P_{T})$
$\overline{B}\to D\overline{T}$, $\Delta S=0$ $\checkmark$ X $\checkmark$ $(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\overline{B}\to\overline{D}T$, $\Delta S=0$ X $\checkmark$ X $(C^{\prime}_{\bar{D}T}+P^{\prime}_{T})$
Before we end this section, it will be useful to compare the participation of $T^{*}_{c\bar{s}0}(2900)^{++}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{cs0}(2870)^{0}$ in $\overline{B}\to D\overline{T}$ and $\overline{B}\to\overline{D}T$ decays in $\Delta S=-1$ and $\Delta S=0$ transitions. We show in Table 11 such a comparison with the relevant topological amplitudes listed. It is evident that using only $\Delta S=-1$ transitions, one is unable to check if these three states belong to the same multiplet or not, as $(C_{\bar{T}}+P_{\bar{T},1})$ and $(C_{\bar{D}T}+P_{T})$ are unrelated. Therefore, to verify that, in any case, we need to consider the CKM suppressed $\Delta S=0$ transitions as well.
@@ V.1.2 Topological amplitudes in $\overline{B}\to TS$ decays
Note that in general, topological diagrams in $\overline{B}\to TS$ decays, see Fig. 7, have more $q\bar{q}$ line(s) produced than those in $\overline{B}\to TP$, see Fig 6. According to the OZI rule, we will expect topological amplitudes in $\overline{B}\to TS$ decays to receive some suppressions. However, since most of the diagrams in $\overline{B}\to TS$ decays are rather different from those in $\overline{B}\to TP$ decays, it is hard to provide a reliable comparison on rates on this basis.
@@ V.2.1 $\overline{B}\to TP$ decay amplitudes
The $\overline{B}\to TP$ decay amplitudes, decomposed in these topological amplitudes, are shown in Table 12 for $\Delta S=0$ and 1-1 transitions in scenario I, and in Tables 13 and 14 for $\Delta S=0$ and $\Delta S=-1$ transitions in scenario II, respectively. These results can be applied to $\overline{B}\to TV$ decays by replacing $P$ with $V$ suitably.
@@
[…]
@@
# Mode $A(\overline{B}_{q}\to T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}P)$ # Mode $A(\overline{B}_{q}\to T_{cq[\bar{q}^{\prime}\bar{q}^{\prime\prime}]}P)$
$2^{*}$ $\overline{B}{}^{0}\to T^{+}_{c\bar{s}}K^{-}$ $\frac{1}{\sqrt{2}}(C_{TP}-E_{\bar{B}T})$ $2^{*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}\pi^{-}$ $\frac{1}{\sqrt{2}}(C_{TP}-C_{PT,1})$